Sorry, you need to enable JavaScript to visit this website.

MCST Banner Image

pm1weekly_2004-290_2004-296

Mission Operations Days: 2004/290 to 2004/296
October 16, 2004 00:00:00 GMT to October 22, 2004 00:00:00 GMT


Aqua Spacecraft and MODIS Instrument Status:

PM-1 "Aqua" is in Fine Pointing Normal Mode
FM1 MODIS is in B-Side Nominal Science Mode
ComponentState(s)Comment
BlackbodyA Off; B On (285K)Nominal Science Mode
Calibration ElectronicsA Off; B OnNominal Science Mode
Control ProcessorA Off; B OnNominal Science Mode
Nadir DoorUnlatched, OpenNominal Science Mode
Space View DoorUnlatched, OpenNominal Science Mode
Solar Diffuser DoorUnlatched, ClosedNominal Science Mode
FDDI FormatterA Off; B OnNominal Science Mode
FDDI PortA Off; B OnNominal Science Mode
FIFO Memory#1 & #2 Off; #3 & #4 OnNominal Science Mode
Format ProcessorA Off; B OnNominal Science Mode
Power Supplies#1 Off, #2 OnNominal Science Mode
PV VIS FPAA Off; B OnNominal Science Mode
PV NIR FPAA Off; B OnNominal Science Mode
PV SMIR FPAA Off; B OnNominal Science Mode
PV LWIR FPAA Off; B OnNominal Science Mode
PC LWIR FPAA Off; B OnNominal Science Mode
RC Outgas HeatersAll 3 (CS, IS, OS) OffNominal Science Mode
RC LWIR FPA HeaterOffNominal Science Mode
RC SMIR FPA HeaterOn (83K)Nominal Science Mode
Scan AssemblyA Off; B OnNominal Science Mode
SDSMOffNominal Science Mode
SRCAOffNominal Science Mode
PS1 Survival HeatersEnabledNominal Science Mode
PS2 Survival HeatersEnabledNominal Science Mode
Timing GeneratorA Off; B OnNominal Science Mode
Flight SoftwareRev BJ + 1 set of patches
  • CP07 to Safe Mode - No FR OFF Command (1 patch)
Nominal Science Mode
Inhibit Ids SetNoneNominal Science Mode

This Week’s Completed MODIS Activities:

Saturday, October 16, 2004

None

Sunday, October 17, 2004

None

Monday, October 18, 2004

None

Tuesday, October 19, 2004

None

Wednesday, October 20, 2004

None

Thursday, October 21, 2004

None

Friday, October 22, 2004

2004/296 18:50:56-18:51:40 Real-Time MOD_MEM_PM_DUMP('motor_constants',35,0,0,1,2)
2004/296 18:51:59-18:52:42 Real-Time MOD_MEM_PM_DUMP('motor_trigger_data',60,0,0,1,2)
2004/296 18:52:59-18:53:28 Real-Time MOD_MEM_PM_DUMP('motor_pos_when_hit',12,0,0,1,1)
2004/296 18:53:40-18:54:14 Real-Time MOD_MEM_PM_DUMP('door_data',12,0,0,1,1)

This Week’s Scheduled MODIS Activities Not Completed:

None

Upcoming MODIS Events:

Calibrations for October include biweekly SDSM, quarterly SRCA Full Spectral, and quarterly Blackbody.

A request by MCST has been made to perform SRCA 10W and 1W continuous mode Radiometric observations. These activities will be coordinated for November. These tests will be similar to the ones performed in October 2003.

Saturday, October 23, 2004

2004/297 21:43:52 MCL MOD_OA03_FR_NIGHT_RATE (03:23 early)
2004/297 22:24:47-22:31:33 MCL MOD_OA08_ROLL_MAN
2004/297 22:40:03 MCL MOD_OA03_FR_DAY_RATE (03:23 late)

Monday, October 25, 2004

2004/299 20:34:38-20:48:36 MCL MOD_OA16_SDSM_SCREEN.2 OA-16: SD/SDSM Screened - B-Side
2004/299 22:13:31-22:27:29 MCL MOD_OA15_SDSM_OPEN.1 OA-15: SD/SDSM Open - B-Side

Tuesday, October 26, 2004

2004/300 00:30:00-00:30:16 MCL MOD_SRCA_SETUP_NO_L2.2 SRCA setup activity using lamp #4 in lieu of lamp #2
2004/300 00:41:33-01:33:15 MCL MOD_OA22_SRCA_FULL_SPECT_1.1 OA-22: SRCA Full Spectral Calibration Part 1
30W-lamp data collection is from 00:57:27-01:33:14
2004/300 02:37:31-03:12:03 MCL MOD_OA22_SRCA_FULL_SPECT_2.1 OA-22: SRCA Full Spectral Calibration Part 2
30W-lamp data collection is from 02:37:31-03:12:02
2004/300 03:50:00-03:50:16 MCL MOD_SRCA_SETUP_NO_L2.2 SRCA setup activity using lamp #4 in lieu of lamp #2
2004/300 03:59:18-04:50:19 MCL MOD_OA22_SRCA_FULL_SPECT_3.1 OA-22: SRCA Full Spectral Calibration Part 3
10W-lamp data collection is from 04:15:45-04:50:18
2004/300 05:59:19-06:37:20 MCL MOD_OA22_SRCA_FULL_SPECT_4.1 OA-22: SRCA Full Spectral Calibration Part 4
10W-lamp data collection is from 05:59:19-06:37:20

Wednesday, October 27, 2004

2004/301 03:00:00 MCL MOD_BB_TEMP_SET_270K_B.1 BB Temp Set to 270K for CP B Coefficients
2004/301 21:00:00 MCL MOD_BB_TEMP_SET_280K_B.1 BB Temp Set to 280K for CP B Coefficients

Thursday, October 28, 2004

2004/302 01:00:00 MCL MOD_BB_TEMP_SET_285K_B.1 BB Temp Set to 285K for CP B Coefficients
2004/302 04:00:00 MCL MOD_BB_TEMP_SET_290K_B.1 BB Temp Set to 290K for CP B Coefficients
2004/302 08:00:00 MCL MOD_BB_TEMP_SET_315K_B.1 BB Temp Set to 315K for CP B Coefficients
2004/302 08:00:30 MCL MOD_BB_DCYCLE_FULL.1 Set blackbody heater duty cycle to FULL
2004/302 12:00:00 MCL MOD_BB_TEMP_SET_270K_B.1 BB Temp Set to 270K for CP B Coefficients
2004/302 12:00:30 MCL MOD_BB_DCYCLE_THIRD.1 Set blackbody heater duty cycle to THIRD

Friday, October 29, 2004

2004/303 08:00:00 MCL MOD_BB_TEMP_SET_285K_B.1 BB Temp Set to 285K for CP B Coefficients

Maneuvers:

The next Aqua MODIS Lunar Calibration Roll will be on DOY 2004/297 (October 23, 2004). The roll angle will be -16.6865 degrees for a nominal 6:30 (390 second) collection.

The Aqua satellite has executed a series of five Inclination Adjust Maneuvers (IAM) between DOY 2004/258 and 2004/317 (September 14, 2004 and November 12, 2004, respectively). The dates for each:

2004/258 (September 14, 2004) IAM #3 Executed: 15:18:11-15:49:31 slewed off nominal pointing 550 second thruster burn
2004/265 (September 21, 2004) IAM #4 Executed: 15:23:36-15:54:56 slewed off nominal pointing 550 second thruster burn
2004/272 (September 28, 2004) IAM #5 Executed: 15:28:59-16:00:19 slewed off nominal pointing 550 second thruster burn
2004/279 (October 5, 2004) IAM #6 Executed: 15:34:22-16:05:42 slewed off nominal pointing 550 second thruster burn
2004/286 (October 12, 2004) IAM #7 Executed: 15:39:46-16:11:06 slewed off nominal pointing 550 second thruster burn

No official date has yet been scheduled for the Aqua Deep Space Calibration pitch maneuver(s).

MODIS Anomalies:

None

General Instrument Comments:

FM1 MODIS is in Science Mode. The SVD and NAD are Open, the survival heaters are enabled, and the Black Body is heated to 285 K.

MODIS Telemetry Trends:

The FM1 MODIS instrument telemetry is trending nominally.

Non-MODIS Significant Events:

The EMOS mission status webpage (which includes instrument data loss records) can be found at the following URL. Note that a password is now required to access the EOS mission links webpages. There are directions at the URL below for how to obtain a username/password.

http://jupiter02.gsfc.nasa.gov/links/links.htm

Limited Life Item Status:

SRCA 10W Lamp #1: 236.5 of 500 hours
SRCA 10W Lamp #2: 184.3 of 500 hours
SRCA 10W Lamp #3: 200.0 of 500 hours
SRCA 10W Lamp #4: 73.2 of 500 hours

SRCA 1W Lamp #1: 508.4 of 5000 hours
SRCA 1W Lamp #2: 269.8 of 5000 hours

Solar Diffuser Door: 2272 of 3022 Movements
Nadir Aperture Door: 1053 of 1316 Movements
Space View Door: 632 of 1316 Movements